OpenStudy (anonymous):

The highest elevation in Minnesota is about 700 meters above sea level. That is 6 2/3<--(mixed number) times higher than the highest elevation in Florida. To the nearest meter, what is the highest elevation in Florida?

OpenStudy (bohotness):

what do you think it is

Directrix (directrix):

@jazelly_ Then, 6 2/3 * F = M where F is the highest elevation in Florida and M is the highest elevation in Minnesota. 6 2/3 * F = 700 20/3 * F = 700 Multiply both sides by 3/20 to solve for F. If you don't like looking at an equation, then divide 700 by 20/3